OpenStudy (anonymous):

2. Find the equation of a line, in slope-intercept form, perpendicular to y = 2x + 3 passing through point (4, - 5).

OpenStudy (anonymous):

m=-1/2 y-(-5)=-(1/2)(x-4) y+5=-(1/2)x+2 y=-(1/2)x-3